OpenStudy (anonymous):

Find the Critical points of the function: f(x)= 2x^3+x^2+2x

OpenStudy (anonymous):

solve it as 2x^3+x^2+2x=0 and find x

OpenStudy (anonymous):

I think i'm supposed to take the derivative of the function first, before I find x.

OpenStudy (anonymous):

do they give you a domain?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay. i know how to do these! i just reviewed them today!

OpenStudy (mimi_x3):

f'(x) = 6x^2 + 2x + 2 Are you finding the Turning Point ? I don't get what critical point means =/

OpenStudy (anonymous):

find the derivative, then see where x is limited. so like what x cannot equal, those are the critical points.

OpenStudy (anonymous):

thats where i am stuck. the derivative is 6x^2+2x+2 but then i don't know how to solve for x here

OpenStudy (anonymous):

I'm just messing up on the simple algebra at the end

OpenStudy (anonymous):

critical points....take the first derivatice and set it to zero which mimi did. just do some subtracting and dividnt and you will get your crit points. Extrema occure around crit pnts

OpenStudy (anonymous):

i don't think there are any critical points

OpenStudy (anonymous):

yes there is

OpenStudy (anonymous):

you have to use quadtratic formula or complete the square to find your roots

OpenStudy (anonymous):

oh okay that explains it, thanks you

OpenStudy (anonymous):

thank you*

OpenStudy (anonymous):

that is a parabola so there is a minima

OpenStudy (mimi_x3):

w8, does the critical point mean the turning point ?

OpenStudy (anonymous):

Also looking at the graph of this... there are no x intercepts so that will save you some trouble

OpenStudy (anonymous):

critical points are where d/dx=0 or d/dx=dne

OpenStudy (anonymous):

if f'(c)=0 or it is not differntiable at c then c is a critical number of f

OpenStudy (anonymous):

oooohh. i got it.

OpenStudy (mimi_x3):

Oh okay, then its the turning point then xD

OpenStudy (anonymous):